Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Sharon Henry, Bobbie Parker, Suzy Shaw, Leota WhitfieldAnnouncements
Ongoing Systems Updates
- New phones in Systems - Origin of phone glitch that occurred this morning is unknown. Functions have returned to normal now. When no one is in room 10 (where the phone will usually ring), the student phone will need to be activated to feature 871 - which will cause the phones to ring in the back area. Upon return, feature #871 will disengage - but only from the student phone unit. The last person out will be responsible for changing the setting - especially after 5 PM. Messages - ‘feature 1’ - basically, a flag that you have a message in your mailbox. Zone 1 is the zone for Systems. We can broadcast to Special Collections & LAC. Intercom works exceptionally well. Different phone ringing tones are available. The front desk phone unit is different, basically it is the operator phone. Has LCD for each individual line. Hold, camp, park - all similar functions. Camp is to be used if the line is busy & the person wants to hold for them.
- The wiring and renovation of Library East is finished. The matter of the fire alarms and door locks is still up in the air.
- Lee has moved out to her room 10 cubicle, Rich is eligible to move into her old office when he is ready.
- A temp worker may be hired to fill in Destry’s position. A server may be set up in Destry’s office.
- Sharon’s new schedule for the summer term while she is enrolled in class: MTWF: 6:30 AM-5:30 PM. Class: MTWF 11:15 AM-12:15 PM. Off Thursdays.
- It is unknown if 1 or 2 extra Cows have been ordered. Betty Mitchell will be consulted.
- Of the Cows received, 20 are on hold for MSL. Only 2 for Special Collections and one 17" for MSL remain. All the rest have been distributed.
- People are wanting to bring back rebuilts (Cows have replaced them). All Macs must come back first.
- Need to make a count of Macs still in use & get them out of service. Dr. Harrer is the exception.
- Rich will deliver the 156K modem for MSL to John Ashcraft.
- Jim Liversidge will be taking over as the liaison for Special Collections.
- Change in meeting time: liaisons will now meet on the first & third Thursdays of the month. The other 2 Thursdays can be used to schedule special training sessions.
- The time and place of the meetings will remain the same for now, although it could eventually change to MSL.
- Recently, there has been more activity on the Liaison list - as long as the discussion is among themselves & not so much directed at Bill for answers - this is good.
Software Projects Status
- Turmoil surrounds public printing - Capitol Copy is pulling out - and they are taking their switch boxes with them. Now the printing has to be set up from a single computer to a single print station. We had to make 2 dozen cables available (with no notice) on Friday.
- Lee found a couple of switches which were ordered for Reference - it will allow 2 computers to print, but no more. It is really not designed to work with the new printers.
- Xerox has the new copier/printing contract. Their system is to be completely networked - no connections to the computers. It will have one control station - computers will be routed to 2 dedicated print servers - 1 in MSL, the other in West (maybe East). This will keep the traffic regional. Maintenance & service responsibility lies with Xerox. Systems will not be the operations for this system - we will provide them with several IP’s. Chris Walton is now a consultant with Xerox.
- 4th, 5th & 6th floors - dumb terminal status is unknown. We’ll probably go with barely loaded micro computers which will act like dumb terminals. There is also no decision on printing from the dumb terminals. We will try to get the Control Units out as soon as possible. We are currently saving $100,000 this year & probably every year here after by moving off the mainframe. It is a very cost effective move.
- Education & Library West are scheduled to go with the switches on May 19. NERDC is nervous about this change. The FIDDI will be disconnected from the router & plugged into the switch. Will have to change the subnet mask - not sure of the effect on the NT server.
- Library East will switch one week after West - everything will have to go down temporarily.
- MSL - Swamp II still pending on furniture.
- ILL migration to MS SQL is in the testing phase.
- Invoicing was done last week - Joan did it with Winston’s supervision. The invoices were printed successfully on a LaserWriter, they were ecstatic. They may not need their 387 printer anymore - definitely won’t need the EPSON.
- Brian presented his demo to Special Collections. They have requested a few changes - it is still ongoing.
- Sharon is back working on the Faculty portion of the Personnel migration. There still seems to be the occasional glitch on the OPS side.
- Bill heard a rumor that David Allerton is finished with Mellon - Winston will try to confirm.
- Patron Update for Summer A & B terms is scheduled.
- Cecilia (Accounting) - her HD died, Winston gave her a tutorial on backing up her files.
- Liaisons have been discussing the use of Public Folders on Exchange - we may use them as a place to drop backups.
- Application Sharing - started tests for individual work stations - with mixed results. Will probably pull it off NT3. Liaisons will track installs - should be done this week.
- No students this week.
- New students will start, Jason returns - on Monday, May 11.
- Bill & Suzy out May 14 - May 15.
- Debra out May 15.
- Brian out May 21-22.
- Memorial Day observed May 25.
- Bobbie out June 8 - 19.
- Debra out half day June 16, all day June 17.
- Bill & Winston out June 15-24.
- Independence Day observed July 3.
- Debra out July 7 - 10.
- Bobbie out July 13 - 17
- Brian out July 16 & 17.
Back to the Systems Homepage
Last updated March 7, 2000
by Debra Harris